Find the dimensions of a right circular cylinder that is open on the top, is closed on the bottom, has volume V , and uses the least amount of material.
1
Expert's answer
2020-04-08T12:09:00-0400

Solution:


Consider the right circular cylinder of radius "r" and height "h".

Given, the volume of the cylinder is "V" , so


"\pi r^2h=V""h=\frac{V}{\pi r^2}"



Total surface area of cylinder(excluding top) is given by,


"A=2\pi rh+\pi r^2"

Substitute "\frac{V}{\pi r^2}" for "h" into equation "A=2\pi rh+\pi r^2" to express the area in terms of single variable "r" as,


"A(r)=2\pi r(\frac{V}{\pi r^2})+\pi r^2""A(r)=\frac{2V}{ r}+\pi r^2"

Differentiate "A(r)" with respect to "r" as,


"A'(r)=\frac{d}{dr}(\frac{2V}{ r}+\pi r^2)=-\frac{2V}{r^2}+2\pi r"

Equate the derivative "A'(r)" to zero to obtain,


"-\frac{2V}{r^2}+2\pi r=0""r=\sqrt[3]{\frac{V}{\pi}}"

Find second derivative to check for maxima/minima as,


"A"(r)=\frac{d}{dr}(-\frac{2V}{r^2}+2\pi r)=\frac{4V}{r^3}+2\pi"



Here, "A"(r)=\frac{4V}{r^3}+2\pi" is positive for any positive value of "r". So, minimum occurs at "r=\sqrt[3]{\frac{V}{\pi}}" .


Now, plug "r=\sqrt[3]{\frac{V}{\pi}}" into relation "h=\frac{V}{\pi r^2}" and simplify for "h" as,



"h=\frac{V}{\pi (\frac{V^\frac{2}{3}}{\pi^\frac{2}{3}})}=\frac{V^\frac{1}{3}}{\pi ^\frac{1}{3}}=\sqrt[3]{\frac{V}{\pi}}"



Therefore, the dimension of right circular cylinder is: Radius"(r)=\sqrt[3]{\frac{V}{\pi}}" and Height"(h)=\sqrt[3]{\frac{V}{\pi}}" .